CH/Agen/96/01

Agenda

Charissa Group Meeting To be held on: Wednesday, 1 May 1996, 11.30 am

To be held at:  Birmingham

Distribution: char-meeting@nnga.daresbury.ac.uk

1. Apologies
2. Minutes of 11 Oct 1996 and 24 Nov 96
3. Matters Arising
4. Update on EPSRC matters
5. Progress on MEGHA hardware tests at ANU  (SPC)
6. Progress on MEGHA read-out tests at Birmingham in July (SJH)
7. Discussion on plans for MEGHA commissioning at ANU
8. Report on recent GANIL run  (WNC)
9. Progress on GANIL array and read-out  (NMC)
10. Report on recent Strasbourg visit  (BRF)
11. Discussion on future ANU running
12. Discussion on future GANIL running
13. Discussion on future Strasbourg running
14. Discussion on organisation of collaboration
15. Progress on papers
16. Any other business
17. Date of next meeting

W.N. Catford (Secretary) 1 May 1996